3. The demand for cookies is given by P = 36.2 – 0.2QD where QD is the quantity demanded of the cookies (number of cookies per day) and P is the price of cookies ($/cookie). Also, the supply of cookies is given by P = 1.4 + 0.3QS where QS is the quantity supplied of the cookies (number of cookies per day) and P is the price of cookies ($/cookie).

Find the equilibrium price and quantity. Show your work. (2)
1
Expert's answer
2020-09-28T10:04:13-0400

At equilibrium, quantity demanded "(""QD)" = quantity supplied "(QS)"

"36.2 - 0.2Q = 1.4 + 0.3Q"

"36.2 - 1.4 = 0.3Q + 0.2Q"

"34.8 = 0.5Q"

"Q = \frac{34.8}{0.5} = 69.6" cookies per day (equilibrium quantity)

"P = 36.2 - 0.2(69.6) = \$22.28"

per cookies which is the equilibrium price